The inside of a baking oven is at a temperature of 177degrees celcius. The wall of an oven is made up of 4mm thick stainless steel of thermal conductivity 15.5W/mk. Determine the heat conduction heat loss per unit area from the oven if outside temperature is at 27degrees celcius above room temperature
1
Expert's answer
2017-01-23T06:50:35-0500
The heat conduction heat loss per unit area from the oven is
Q=kA ΔT/Δx
Q/A=k ΔT/Δx=15.5 (177-27)/0.004=581250 W/m^2 .
Answer: 581250 W/m^2 .